OpenStudy (anonymous):

(2/x+2)+(3/x)=(-x/x+2) multiply each side by LCD which is x(x+2) answer shld be -3 but idk how to get to it pls show work

OpenStudy (anonymous):

\[(2/x+2) \times ({x/x+2}) + (3/x) \times (x/x+2) = -(x/x+2) \times (x/x+2)\] take the LCM and solve it up a little and you will end up with \[x^2 + 5x + 6 = 0\] and just solve the quadratic. You will get two answers, -3 and -2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

of u mean x= -b + or - square root blah blah blah thing right?

OpenStudy (anonymous):

yeah theyre right